Fermi National Laboratory


Accelerator Update

Monday July 26
The day shift began with Operations monitoring store 3674, with the Antiproton Source (Pbar) stacking, and with MiniBooNE and the Meson MCenter experiments taking beam.

Experts for the Recycler and Main Injector (MI) began a Time Line Generator (TLG) study at 9:33 AM.

A power supply expert investigated the MiniBooNE Horn power supply problem from over the weekend at 10:11 AM. The Horn was off for about five minutes.

At 10:15 AM, Operations reported that a Booster RF station (BRF4) was tripping off repeatedly. An expert investigated.

Operations set up Pbar to conduct reverse proton studies at 10:27 AM.

A RF expert reported at 10:57 AM that he had repaired MIRF10 & 11, and he was checking on stations 13-16.

Operations turned off the Switchyard (SY) SEPTA at 11:01 AM in order to continue with the Pbar reverse protons studies.

Tevatron (TeV) experts began end-of-store studies at 12:04 PM.

Operations reported at 12:27 PM that the MiniBooNE Horn was back on and running, but at a lower current.

BRF4 tripped off at 12:30 PM. It reset okay.

Cryo system experts were given permission at 1:07 PM to work on the TeV sectors A3 and F4 wet engines after the store ends.

Operations terminated store 3674 at 1:09 PM.

Operations made the Collider Detector at Fermilab (CDF) and D-Zero (D0) experiments ready for access at 1:34 PM.

CDF completed their access at 2:45 PM.

BRF4 tripped off at 2:52 PM. It reset okay.

Pbar experts completed their reverse proton studies at 3:15 PM. Stacking resumed about ten minutes later.

MIRF8 tripped off at 3:26 PM. It reset okay.

D0 completed their access about 3:30 PM. Operations began the TeV turn on.

Operations put the TeV through a dry squeeze at 4:20 PM.

Operations put the TeV into shot setup at 4:57 PM. (Operations disables all SY beam during every shot setup.)

A TeV RF station (TRF8) tripped off at 7:34 PM on a heater ground fault indication. It was left off.

Operations established store 3676 at 7:38 PM with an initial luminosity of 72.41E30. Pbar resumed stacking.

Operations reestablished beam to MCenter at 7:52 PM.

Tuesday July 27
The midnight shift began with Operations monitoring store 3676, with Pbar stacking, and with MiniBooNE and MCenter taking beam.

At 1:01 AM, a Booster vacuum pump (IPRF17) tripped off and took down BRF7 & 8. Everything reset okay.

At 1:25 AM, a Linac RF station (LRF3) tripped off. At 2:26 AM, LRF5 tripped off and at 5:37 AM KRF1 tripped off. They all reset okay.

At 5:41 AM, a MI RF station (MIRF15) tripped off and would only come partially back on. Operators bypassed the station.

A Recycler expert began a Time Line Guide (TLG) study at 6:30 AM.

At 6:55 AM, the MiniBooNE Horn tripped off for the first of many times in thirty minutes. Operations reduced the rep rate and this resolved the trips.

An RF expert repaired MIRF15. It was online and back in the grid by 9 AM.

At 2:02 PM, MIRF17 tripped off. It reset okay.

At 2:28 PM, MIRF18 and BRF17 tripped off. MIRF18 reset okay, but operators had trouble resetting BRF17. They bypassed the station and called experts.

RF expert reported at 3:10 PM that BRF15 & 17 needed a tunnel access to repair.

MI slipstacking studies ended at 4:35 PM.

At 4:52 PM, BRF7 tripped off. It reset okay.

NuMI experts began a brief power supply test at 5:12 PM.

A MI expert began a damper study at 5:35 PM.

A TeV expert began an end-of-store study at 7:05 PM.

Operations terminated store 3676 at 8:10 PM.

Operations prepared D0 for an access at 8:23 PM. They came out of access an hour later.

An RF expert came in a brought up TRF8 at 9:31 PM.

Operations put the TeV into shot setup at 10:14 PM.

A Booster power supply (EAPS) tripped off at 11:15 PM due to a front panel breaker trip on BRF4. It reset okay.

At 11:51 PM, LRF5 tripped off. It reset itself okay.

Wednesday July 28
The midnight shift began with the TeV in shot setup, and with MiniBooNE taking beam.

Operations established store 3678 at 12:28 AM with an initial luminosity of 84.53E30. Pbar resumed stacking.

Operations resumed sending beam to MCenter at 12:49 AM.

At 1:20 AM, LRF5 tripped off. It tripped off a few more times while operators tried to reset it. Eight minutes later, it tripped off again.

At 2:11 AM, KRF1 tripped off and wouldn't reset. Operator called in a Linac technician.

An operator noticed at 3:06 AM that a possible problem with the KRF1 LCW flow meter was keeping the station off. He grabbed a wrench and convinced the meter to resume working. The station then came on and resumed normal operations.

Operations halted stacking at 6 AM. The Recycler stashed antiprotons from Pbar.

Operations began preparing the Linac, Booster and Pbar tunnels for access at 7 AM.

Plans for Wednesday
The Linac, Booster, and Pbar will be in access for three to four hours this morning. Pbar will begin stacking this afternoon. Operations will maintain the store. The next shot is expected sometime on Thursday after a stack has been built up.

Accelerator Update Archive

More Information
For Tevatron luminosity charts and the current status of Fermilab's accelerators and detectors (live!), please go to Fermilab Now

Comments and Suggestions
What do you think about the Accelerator Updates? Please send comments and suggestions to: accelupdates@fnal.gov.



last modified 7/28/2004   email Fermilab